# Compatible Poisson Structures On Multiplicative Quiver Varieties

### v2

@@ Abstract.
Any multiplicative quiver variety is endowed with a Poisson structure constructed by Van den Bergh through reduction from a Hamiltonian quasi-Poisson structure. The smooth locus carries a corresponding symplectic form defined by Yamakama through quasi-Hamiltonian reduction. In this note, we include the Poisson structure as part of a pencil of compatible Poisson structures on the multiplicative quiver variety. The pencil is defined by reduction from a pencil of Hamiltonian quasi-Poisson structures which has dimension $\ell(\ell-1)/2$, where $\ell$ is the number of arrows in the underlying quiver. For each element of the pencil, we exhibit the corresponding compatible symplectic or quasi-Hamiltonian structure. We comment on analogous constructions for character varieties and quiver varieties. This formalism is applied to the spin Ruijsenaars-Schneider phase space in order to explain the compatibility of two Poisson structures that have recently appeared in the literature.
@@ 1. Introduction
Acknowledgments. It is my pleasure to thank Oleg Chalykh, Laszlo Fehér, David Fernández, Anne Moreau and Olivier Schiffmann for stimulating discussions. I am especially indebted to Laszlo Fehér for comments on a first version of this paper, and to Anton Alekseev for encouraging me to develop Remark 3.21 and Section 4. This project received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101034255, while affiliated to Laboratoire de Mathématiques d’Orsay.
@@ 2.2.2. Definitions and standard results
A $G$-invariant bivector $P$ on $M$ is quasi-Poisson if $[P,P]=\phi_{M}$, where [,][-,-][ - , - ] denotes the Schouten-Nijenhuis bracket on $M$. Then $(M,P)$ is called a quasi-Poisson variety.
@@
Fix a $G$-invariant $2$-form $\omega$ on $M$ and a $G$-equivariant morphism $\Phi:M\to G$. We say that the triple $(M,\omega,\Phi)$ is a quasi-Hamiltonian variety (with moment map given by $\Phi$) if the following three conditions are satisfied:
@@
There is a $1$-$1$ correspondence between structures of non-degenerate Hamiltonian quasi-Poisson variety and those of quasi-Hamiltonian variety on $M$ admitting the (same) moment map $\Phi:M\to G$ for which the associated quasi-Poisson bivector $P$ and $2$-form $\omega$ are related by:
@@ 2.3.1. Quiver representations
Fix a quiver $Q$ and a dimension vector $\mathbf{n}=(n_{s})\in\mathbb{Z}_{\geq 0}^{I}$. A representation of dimension $\mathbf{n}$ of the path algebra $\mathbb{C}Q$ is an element $\rho\in\operatorname{Hom}_{B}(\mathbb{C}Q,\operatorname{End}(\mathbb{C}^{\mathbf{n}}))$ such that, for each $s\in I$, $\rho(e_{s})$ is the projection on the summand $\mathbb{C}^{n_{s}}$ of $\mathbb{C}^{\mathbf{n}}:=\oplus_{r\in S}\mathbb{C}^{n_{r}}$. Using the convention of § 2.1, we see that the matrix $\rho(a)$, $a\in Q$, may only have a nonzero block of size $n_{t(a)}\times n_{h(a)}$ corresponding to a linear map [^3] We warn the reader that we use Van den Bergh’s convention [39, 40] for quiver representations, which is the opposite of the convention taken in [14, 42]. $\mathbb{C}^{n_{h(a)}}\to\mathbb{C}^{n_{t(a)}}$. Therefore the space of all representations of dimension $\mathbf{n}=(n_{s})\in\mathbb{Z}_{\geq 0}^{I}$, denoted $M_{Q}:=M_{Q}(\mathbf{n})$, is an affine space $\mathbb{A}^{d}_{\mathbb{C}}$ of dimension $d:=\sum_{a\in Q}n_{t(a)}n_{h(a)}$. This turns $M_{Q}$ into an affine variety. Rather than working with points as representations of $\mathbb{C}Q$, we view a point $\mathtt{X}\in M_{Q}$ as being parametrized as $\mathtt{X}=(\mathtt{X}_{a})_{a\in Q}$ with $\mathtt{X}_{a}:=\rho(a)$ for $\rho$ the corresponding representation.
@@
[…]
@@ 2.3.2. Definition
Introduce the smooth open affine subvariety $M_{\overline{Q}}^{\circ}:=\{\mathtt{X}\mid\det(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})\neq 0\,\forall a\in\overline{Q}\}\subset M_{\overline{Q}}$. Fix a total ordering $<$ on $\overline{Q}$. We reproduce from (1.1) the $\operatorname{GL}_{\mathbf{n}}$-equivariant morphism
@@
where factors are taken from left to right with respect to the ordering $<$ on the arrows. Similarly, set $\Phi_{a}:=\prod_{b<a}(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{\epsilon(b)}$ for all $a\in\overline{Q}$. Finally, fix $\mathbf{q}:=(q_{s})\in(\mathbb{C}^{\times})^{I}$ and denote in the same way the central element $(q_{s}\operatorname{Id}_{n_{s}})_{s}\in\operatorname{GL}_{\mathbf{n}}$. Note that $\Phi^{-1}(\mathbf{q})$ is a $\operatorname{GL}_{\mathbf{n}}$-invariant closed subvariety of $M_{\overline{Q}}^{\circ}$, which is empty if $\mathbf{q}^{\mathbf{n}}:=\prod_{s\in I}q_{s}^{n_{s}}\neq 1$ because $\det\circ\,\Phi=1$. This subvariety parametrizes representation of the multiplicative preprojective algebra $\Lambda^{\mathbf{q}}(Q)$, cf. [14]. We are in position to define the main objects at stake, see [42, §2.2] for the relevant geometric invariant theory of quivers.
@@
$P:=$ $\,\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})\,\partial_{a}\wedge\partial_{a^{\ast}}\right]$ (2.15)
@@
$\omega:=$ $\,-\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}\,\mathrm{d}\mathtt{X}_{a}\wedge\mathrm{d}\mathtt{X}_{a^{\ast}}\right]$ (2.16)
@@ 3.1.1. First properties
    $(P_{j})_{j\in J}$ defines a quasi-Poisson pencil of order $r$ centered at $P_{0}$ if and only if $(P_{j})_{j\in J}$ defines a $G$-invariant Poisson pencil of order $r$ and, for any $j\in J$, $P_{j}$ satisfies $[P_{0},P_{j}]=0$.
@@
$[P(\underline{z}),P(\underline{z})]=[P_{0}+P(\underline{z}),P_{0}+P(\underline{z})]-2[P_{0},P(\underline{z})]-[P_{0},P_{0}]=-2[P_{0},P(\underline{z})]\,,$ (3.1)
since $P_{0}$ and $P_{0}+P(\underline{z})$ are quasi-Poisson. Fix $j\in J$, and take $z_{k}=0$ for all $k\neq j$ so that (3.1) yields $z_{j}^{2}[P_{j},P_{j}]=-2z_{j}[P_{0},P_{j}]$; varying $z_{j}\in\mathbb{C}$ implies that $[P_{0},P_{j}]=0$. Thus (3.1) is identically zero and each bivector $P(\underline{z})$ is Poisson. By definition, quasi-Poisson bivectors are $G$-invariant, so that $P_{j}=(P_{0}+P_{j})-P_{0}$ is itself $G$-invariant for any $j\in J$. The fact that the order is the same is direct from the definition. The converse is proved similarly.
@@
Finally, we make some observations regarding quasi-Poisson reduction. Given a quasi-Poisson pencil as in Definition 3.1, the family $(P_{0},P_{j})_{j\in J}$ defines a Poisson pencil on $M/\!\!/G$ of order [^5] By abusing terminology, we allow the order to be zero; this occurs e.g. if $M/\!\!/G$ is 00-dimensional. $r^{\prime}\leq r+1$. In the Hamiltonian case, the family $(P_{0},P_{j})_{j\in J}$ defines a Poisson pencil of order $r^{\prime}\leq r+1$ on any reduction $\Phi^{-1}(\mathcal{C})/\!\!/G$ (where $\mathcal{C}$ is the closure of a conjugacy class in $G$); this is also true in the presence of a stability parameter.
@@ 3.1.2. Construction using an abelian group action
Since $P$ is $H$-invariant and $H$ is abelian, we obtain $[\chi_{M},P]=0$ and $[\chi_{M},\chi^{\prime}_{M}]=0$ for any $\chi,\chi^{\prime}\in\mathfrak{h}$. Thus $[P+\psi_{\underline{z}},P+\psi_{\underline{z}}]=[P,P]$ and the bivector is quasi-Poisson.
@@
Under the assumptions of Lemma 3.4, assume that there exists a closed invariant $2$-form $\varpi_{\underline{z}}$ such that $\varpi_{\underline{z}}^{\flat}(\xi_{M})=0$ for all $\xi\in\mathfrak{g}$, and the compatibility condition (2.7) holds for $P+\psi_{\underline{z}}$ with $\omega+\varpi_{\underline{z}}$. Then $(M,P+\psi_{\underline{z}},\Phi)$ is a non-degenerate Hamiltonian quasi-Poisson $G$-variety and $(M,\omega+\varpi_{\underline{z}},\Phi)$ is the corresponding quasi-Hamiltonian $G$-variety.
@@ 3.1.3. The trick: embeddings of $\mathbb{C}^{\times}$ in the center of $G$
Write $1_{\mathfrak{g}_{j}}$ for the image of $1\in\mathbb{C}$ obtained by differentiating the embedding ×$\mathbb{C}^{\times}\hookrightarrow G_{j}$. By definition of the vector field $\mathrm{Inf}^{(j)}$ and (2.6b), we get
@@
(3.6)$_{LHS}=\sum_{i<j}z_{ij}\,(\langle X,\operatorname{tr}(\Phi_{j}^{-1}\,\mathrm{d}\Phi_{j})\rangle\,\mathrm{Inf}^{(i)}-\langle X,\operatorname{tr}(\Phi_{i}^{-1}\,\mathrm{d}\Phi_{i})\rangle\,\mathrm{Inf}^{(j)})\,.$ ) . (3.8)
@@
$-\sum_{i<j}z_{ij}\,\Big{(}\langle X,\operatorname{tr}(\Phi_{i}^{-1}\,\mathrm{d}\Phi_{i})\rangle\,P^{\sharp}(\operatorname{tr}(\Phi_{j}^{-1}\,\mathrm{d}\Phi_{j}))-\langle X,\operatorname{tr}(\Phi_{j}^{-1}\,\mathrm{d}\Phi_{j})\rangle\,P^{\sharp}(\operatorname{tr}(\Phi_{i}^{-1}\,\mathrm{d}\Phi_{i}))\Big{)}\,.$ (3.9)
@@ 3.2. Case of multiplicative quiver varieties
Note that the action $\mathcal{A}_{b}$ (3.11) of $\mathbb{C}^{\times}$ on $M_{\overline{Q}}^{\circ}$ is obtained using the embedding ×$\mathbb{C}^{\times}\hookrightarrow\operatorname{GL}_{n_{t(b)}}$, $\lambda\mapsto\lambda\operatorname{Id}_{n_{t(b)}}$, from the corresponding action considered in Remark 2.11 before fusion. Therefore it suffices to apply Proposition 3.6 to $M_{\overline{Q}}^{\circ}$ before fusion, then perform fusion as in the original works of Van den Bergh and Yamakawa [39, 40, 42]; we conclude by Lemma 3.2. ∎
@@
[…]

### v3

@@ Abstract.
Any multiplicative quiver variety is endowed with a Poisson structure constructed by Van den Bergh through reduction from a Hamiltonian quasi-Poisson structure. The smooth locus carries a corresponding symplectic form defined by Yamakawa through quasi-Hamiltonian reduction. In this note, we include the Poisson structure as part of a pencil of compatible Poisson structures on the multiplicative quiver variety. The pencil is defined by reduction from a pencil of Hamiltonian quasi-Poisson structures which has dimension $\ell(\ell-1)/2$, where $\ell$ is the number of arrows in the underlying quiver. For each element of the pencil, we exhibit the corresponding compatible symplectic or quasi-Hamiltonian structure. We comment on analogous constructions for character varieties and quiver varieties. This formalism is applied to the spin Ruijsenaars-Schneider phase space in order to explain the compatibility of two Poisson structures that have recently appeared in the literature.
@@ 1. Introduction
Acknowledgments. It is my pleasure to thank Oleg Chalykh, Laszlo Fehér, David Fernández, Anne Moreau and Olivier Schiffmann for stimulating discussions. I am especially indebted to Laszlo Fehér for comments on a first version of this paper, and to Anton Alekseev for encouraging me to develop Remark 3.21 and Section 4. This project received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101034255, while affiliated to Laboratoire de Mathématiques d’Orsay. It was completed at IMB which is supported by the EIPHI Graduate School (contract ANR-17-EURE-0002).
@@ 2.2.2. Definitions and standard results
A $G$-invariant bivector $P$ on $M$ is quasi-Poisson if $P,P]=\phi_{M}$, where [,][-,-] denotes the Schouten-Nijenhuis bracket on $M$. Then $(M,P)$ is called a quasi-Poisson variety.
@@
Fix a $G$-invariant 22-form $\omega$ on $M$ and a $G$-equivariant morphism $\Phi:M\to G$. We say that the triple $(M,\omega,\Phi)$ is a quasi-Hamiltonian variety (with moment map given by $\Phi$) if the following three conditions are satisfied:
@@
There is a 11-11 correspondence between structures of non-degenerate Hamiltonian quasi-Poisson variety and those of quasi-Hamiltonian variety on $M$ admitting the (same) moment map $\Phi:M\to G$ for which the associated quasi-Poisson bivector $P$ and $2$-form $\omega$ are related by:
@@ 2.3.1. Quiver representations
Fix a quiver $Q$ and a dimension vector $\mathbf{n}=(n_{s})\in\mathbb{Z}_{\geq 0}^{I}$. A representation of dimension $\mathbf{n}$ of the path algebra $\mathbb{C}Q$ is an element $\rho\in\operatorname{Hom}_{B}(\mathbb{C}Q,\operatorname{End}(\mathbb{C}^{\mathbf{n}}))$ such that, for each $s\in I$, $\rho(e_{s})$ is the projection on the summand $\mathbb{C}^{n_{s}}$ of $\mathbb{C}^{\mathbf{n}}:=\oplus_{r\in I}\mathbb{C}^{n_{r}}$. Using the convention of § 2.1, we see that the matrix $\rho(a)$, $a\in Q$, may only have a nonzero block of size $n_{t(a)}\times n_{h(a)}$ corresponding to a linear map [^3] We warn the reader that we use Van den Bergh’s convention [39, 40] for quiver representations, which is the opposite of the convention taken in [14, 42]. $\mathbb{C}^{n_{h(a)}}\to\mathbb{C}^{n_{t(a)}}$. Therefore the space of all representations of dimension $\mathbf{n}=(n_{s})\in\mathbb{Z}_{\geq 0}^{I}$, denoted $M_{Q}:=M_{Q}(\mathbf{n})$, is an affine space $\mathbb{A}^{d}_{\mathbb{C}}$ of dimension $d:=\sum_{a\in Q}n_{t(a)}n_{h(a)}$. This turns $M_{Q}$ into an affine variety. Rather than working with points as representations of $\mathbb{C}Q$, we view a point $\mathtt{X}\in M_{Q}$ as being parametrized as $\mathtt{X}=(\mathtt{X}_{a})_{a\in Q}$ with $\mathtt{X}_{a}:=\rho(a)$ for $\rho$ the corresponding representation.
@@
$M_{Q}^{\theta.s}=\{\mathtt{X}\in M_{Q}\mid\mathtt{X}\text{ is }\theta\text{-stable}\}\,,\qquad M_{Q}^{\theta.ss}=\{\mathtt{X}\in M_{Q}\mid\mathtt{X}\text{ is }\theta\text{-semistable}\}\,.$ (2.11)
@@ 2.3.2. Definition
Introduce the smooth open affine subvariety $M_{\overline{Q}}^{\circ}:=\{\mathtt{X}\mid\det(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})\neq 0\,\forall a\in\overline{Q}\}\subset M_{\overline{Q}}$. Fix a total ordering << on $\overline{Q}$. We reproduce from (1.1) the $\operatorname{GL}_{\mathbf{n}}$-equivariant morphism
@@
where factors are taken from left to right with respect to the ordering << on the arrows. Similarly, set $\Phi_{a}:=\prod_{b<a}(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{\epsilon(b)}$ for all $a\in\overline{Q}$. Finally, fix $\mathbf{q}:=(q_{s})\in(\mathbb{C}^{\times})^{I}$ and denote in the same way the central element $(q_{s}\operatorname{Id}_{n_{s}})_{s}\in\operatorname{GL}_{\mathbf{n}}$. Note that $\Phi^{-1}(\mathbf{q})$ is a $\operatorname{GL}_{\mathbf{n}}$-invariant closed subvariety of $M_{\overline{Q}}^{\circ}$, which is empty if $\mathbf{q}^{\mathbf{n}}:=\prod_{s\in I}q_{s}^{n_{s}}\neq 1$ because $\det\circ\,\Phi=1$. This subvariety parametrizes representation of the multiplicative preprojective algebra $\Lambda^{\mathbf{q}}(Q)$, cf. [14]. We are in position to define the main objects at stake, see [42, §2.2] for the relevant geometric invariant theory of quivers.
@@
$P$= $\,\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})\,\partial_{a}\wedge\partial_{a^{\ast}}\right]$ (2.15)
@@
$\omega$= $\,-\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}\,\mathrm{d}\mathtt{X}_{a}\wedge\mathrm{d}\mathtt{X}_{a^{\ast}}\right]$ (2.16)
@@ 3.1.1. First properties
    $(P_{j})_{j\in J}$ defines a quasi-Poisson pencil of order $r$ centered at $P_{0}$ if and only if $(P_{j})_{j\in J}$ defines a $G$-invariant Poisson pencil of order $r$ and, for any $j\in J$, $P_{j}$ satisfies $P_{0},P_{j}]=0$.
@@
$P(\underline{z}),P(\underline{z})]=[P_{0}+P(\underline{z}),P_{0}+P(\underline{z})]-2[P_{0},P(\underline{z})]-[P_{0},P_{0}]=-2[P_{0},P(\underline{z})]\,$, (3.1)
since $P_{0}$ and $P_{0}+P(\underline{z})$ are quasi-Poisson. Fix $j\in J$, and take $z_{k}=0$ for all $k\neq j$ so that (3.1) yields $z_{j}^{2}[P_{j},P_{j}]=-2z_{j}[P_{0},P_{j}]$; varying $z_{j}\in\mathbb{C}$ implies that $P_{0},P_{j}]=0$. Thus (3.1) is identically zero and each bivector $P(\underline{z})$ is Poisson. By definition, quasi-Poisson bivectors are $G$-invariant, so that $P_{j}=(P_{0}+P_{j})-P_{0}$ is itself $G$-invariant for any $j\in J$. The fact that the order is the same is direct from the definition. The converse is proved similarly.
@@
Finally, we make some observations regarding quasi-Poisson reduction. Given a quasi-Poisson pencil as in Definition 3.1, the family $(P_{0},P_{j})_{j\in J}$ defines a Poisson pencil on M//GM/\!\!/G of order [^5] By abusing terminology, we allow the order to be zero; this occurs e.g. if M//GM/\!\!/G is 0-dimensional. $r^{\prime}\leq r+1$. In the Hamiltonian case, the family $(P_{0},P_{j})_{j\in J}$ defines a Poisson pencil of order $r^{\prime}\leq r+1$ on any reduction $\Phi^{-1}(\mathcal{C})/\!\!/G$ (where $\mathcal{C}$ is the closure of a conjugacy class in $G$); this is also true in the presence of a stability parameter.
@@ 3.1.2. Construction using an abelian group action
Since $P$ is $H$-invariant and $H$ is abelian, we obtain $\chi_{M},P]=0$ and $\chi_{M},\chi^{\prime}_{M}]=0$ for any $\chi,\chi^{\prime}\in\mathfrak{h}$. Thus $P+\psi_{\underline{z}},P+\psi_{\underline{z}}]=[P,P]$ and the bivector is quasi-Poisson.
@@
Under the assumptions of Lemma 3.4, assume that there exists a closed invariant 22-form $\varpi_{\underline{z}}$ such that $\varpi_{\underline{z}}^{\flat}(\xi_{M})=0$ for all $\xi\in\mathfrak{g}$, and the compatibility condition (2.7) holds for $P+\psi_{\underline{z}}$ with $\omega+\varpi_{\underline{z}}$. Then $(M,P+\psi_{\underline{z}},\Phi)$ is a non-degenerate Hamiltonian quasi-Poisson $G$-variety and $(M,\omega+\varpi_{\underline{z}},\Phi)$ is the corresponding quasi-Hamiltonian $G$-variety.
@@ 3.1.3. The trick: embeddings of $\mathbb{C}^{\times}$ in the center of $G$
Write $1_{\mathfrak{g}_{j}}$ for the image of $1\in\mathbb{C}$ obtained by differentiating the embedding $\mathbb{C}^{\times}\hookrightarrow G_{j}$. By definition of the vector field $\mathrm{Inf}^{(j)}$ and (2.6b), we get
@@
[…]
@@
$\sum_{i<j}z_{ij}\,\Big(\langle X,\operatorname{tr}(\Phi_{i}^{-1}\,\mathrm{d}\Phi_{i})\rangle\,P^{\sharp}(\operatorname{tr}(\Phi_{j}^{-1}\,\mathrm{d}\Phi_{j}))-\langle X,\operatorname{tr}(\Phi_{j}^{-1}\,\mathrm{d}\Phi_{j})\rangle\,P^{\sharp}(\operatorname{tr}(\Phi_{i}^{-1}\,\mathrm{d}\Phi_{i}))\Big)\,.$ (3.9)
@@ 3.2. Case of multiplicative quiver varieties
Note that the action $\mathcal{A}_{b}$ (3.11) of $\mathbb{C}^{\times}$ on $M_{\overline{Q}}^{\circ}$ is obtained using the embedding $\mathbb{C}^{\times}\hookrightarrow\operatorname{GL}_{n_{t(b)}}$, $\lambda\mapsto\lambda\operatorname{Id}_{n_{t(b)}}$, from the corresponding action considered in Remark 2.11 before fusion. Therefore it suffices to apply Proposition 3.6 to $M_{\overline{Q}}^{\circ}$ before fusion, then perform fusion as in the original works of Van den Bergh and Yamakawa [39, 40, 42]; we conclude by Lemma 3.2. ∎
@@
Computing the order $r$ of the pencil inherited by $\mathcal{M}_{\overline{Q},\mathbf{q},\theta}$ appears to be a challenging problem. If the quiver $Q$ contains loops or multiple arrows between two vertices, we expect that $r>1$ (assuming that $\dim\mathcal{M}_{\overline{Q},\mathbf{q},\theta}>0$). We refer to Corollary 5.4 with $d>2$ for a family of quivers where this holds. However, the added bivectors $\psi_{\underline{z}}$ (3.13) may vanish after quasi-Poisson reduction, leading to trivial cases; this occurs for quivers of $A_{l}$ type, for example. Let us see this in greater generalities for a star-shaped quiver. Following e.g. [42, §4], given $k\geq 1$ and $l:=(l_{1},\ldots,l_{k})\in\mathbb{Z}_{\geq 1}^{k}$, the quiver $Q_{k;l}$ corresponding to these data has for vertex set $I=\{0\}\cup\{(k^{\prime},l^{\prime})\mid 1\leq l^{\prime}\leq l_{k^{\prime}},\,\,1\leq k^{\prime}\leq k\}$ and exactly one arrow $(k^{\prime},1)\to 0$ for each $1\leq k^{\prime}\leq k$, and one arrow $(k^{\prime},l^{\prime})\to(k^{\prime},l^{\prime}-1)$ for each $2\leq l^{\prime}\leq l_{k^{\prime}}$ and $k^{\prime}$. We call a quiver star-shaped if it is of the form $Q_{k;l}$ (for some $k,l$), up to changing the orientation of some arrows.